In a recent sample of 85 used cars sales costs, the sample mean was $6,925 with a standard deviation of $3,159. Assume the underlying distribution is approximately normal.
a. what is the random x variable
b.Construct a 95% confidence interval for the population mean cost of a used car.
(i) State the confidence interval. (Round your answers to one decimal place.)
sketch the graph a/2= CL=
Calucate error bound
1
Expert's answer
2020-09-15T17:27:00-0400

Let  the random variable "X" be the cost of a used car.

The provided sample mean is "\bar{X}=6925" and the sample standard deviation is "s=3159." The size of the sample is "n=85" and the required confidence level is 95%.

The number of degrees of freedom are "df=85-1=84," and the significance level is "\alpha=0.05."

Based on the provided information, the critical t-value for "\alpha=0.05" and "df=84"

degrees of freedom is "t_c=1.9886." The 95% confidence for the population "\mu"  is computed using the following expression


"CI=(\bar{X}-\dfrac{t_c\times s}{\sqrt{n}}, \bar{X}+\dfrac{t_c\times s}{\sqrt{n}})="

"=(6925-\dfrac{1.9886\times 3159}{\sqrt{85}}, 6925+\dfrac{1.9886\times 3159}{\sqrt{85}})="

"=(6243.62, 7606.38)"

"EBM=\dfrac{t_c\times s}{\sqrt{n}}=\dfrac{1.9886\times 3159}{\sqrt{85}}=681.38"